# Enhanced pedestal transport driven by edge collisionality on Alcator C-Mod and its role in regulating H-mode pedestal gradients

### v1

@@ 2.4 Inference of neutral density, ionization source, and particle flux
With $n_{e}$, $T_{e}$, and $\epsilon_{\mathrm{Ly}_{\alpha}}$ all mapped to the same radial grid at the midplane, we make use of a collisional-radiative model to infer $n_{0}$, $S_{\mathrm{ion}}$, and $\Gamma_{D}$. A similar procedure has been employed previously, on C-Mod with Dα measurements [14], and on both C-Mod and DIII-D with Lyα [19, 23, 36, 37, 38]. Equations 2-4 summarize the procedure for calculating these quantities.
@@
$\mathcal{P}$ is the photon emissivity coefficient, $S_{H}=\frac{\mathcal{S}}{\mathcal{P}}$ is the ratio between $\mathcal{S}$, the effective ionization coefficient and $\mathcal{P}$, the photon emissivity coefficient. Neutral and ionized molecular contributions are neglibible for Lyα emission. These expressions highlight the implicit dependence of $S_{\mathrm{ion}}$ (and hence $\Gamma_{D,\perp}$) on $n_{e}$ and $T_{e}$.
Figure 3: Typical kinetic profiles as measured by ETS, showing $n_{e}$ (top) and $T_{e}$ (bottom). Profiles in purple are at higher collisionality (higher $n_{e}$ and lower $T_{e}$), the relevance of which is explained in Section 4. Profiles in green are at lower collisionality (lower $n_{e}$ and higher $T_{e}$). The dotted vertical line in gray denotes the separatrix, located at $\psi_{n}=1.0$.
Figure 4: Typical profiles resulting from the analysis described in 2.4, measured inferred from a combination of ETS and LYMID measurements. Profiles span the plasma edge, from the top of the pedestal into the SOL. The bottom two panels show $\Gamma_{D}$ and $D_{\mathrm{eff}}$, which are only valid until $\psi_{n}=1$. The first panel, $n_{0}$ is plotted in a logarithmic scale to show its large variation. As with Figure 3, purple curves correspond to plasmas at higher collisionality than green curves.
Table 1: Ranges of values at the location of the top of the $n_{e}$ pedestal and at the separatrix for the discharges analyzed. $n_{e}$ and $T_{e}$ vary substantially at the separatrix, but primarily $T_{e}$ varies at the pedestal top. In contrast, all neutral quantities vary substantially across discharges and at both the pedestal and the separatrix.
@@
$n_{0}$ (1016 m-3) 0.7 - 2.3 1.5 - 3.8
$S_{\mathrm{ion}}$ (1022 m-3s-1) 4.5 – 16 5.0 – 17
$\Gamma_{D}$ (1020 m-2s-1) 1.4 – 14 3.3 – 17
$D_{\mathrm{eff}}$ (10-2 m2s-1) 2.4 – 13 1.6 – 11
@@
Equations 2 and 3 use a number of simplifying assumptions about the relevant atomic processes, which are outlined in [37]. For Equation 4, we ensure a stationary plasma density time window and assume no poloidal asymmetries in $S_{\mathrm{ion}}$, allowing computation of the cross-field particle flux, $\Gamma_{D,\perp}$ by integrating Equation 3 in a slab geometry. For the remainder of the paper, we drop the $\perp$ subscript and express the cross-field particle flux as $\Gamma_{D}$. The assumption that there is negligible poloidal asymmetry in $S_{\mathrm{ion}}$ is likely the strongest simplifying assumption. Using two distinct Lyα views, an in-out poloidal asymmetry has been observed in the ionization source on DIII-D [37, 39]. LYMID does not distinctly view the inboard and outboard side, and an in-out source asymmetry may well be present on C-Mod as well. Regardless, past work has shown that C-Mod is dominated by main-chamber fueling [40, 23]. This lends confidence to using neutral inferences at the midplane to make conclusions about fueling and transport in the C-Mod pedestal. We test this assertion with EIRENE, a 3D neutral simulation code, in Section 5 and observe good agreement with previous conclusions about main chamber fueling.
@@
Figures 3 and 4 show typical profiles produced with this workflow. Figure 3 shows $n_{e}$ and $T_{e}$ computed from ETS. Figure 4 shows $n_{0}$, $S_{\mathrm{ion}}$, $\Gamma_{D}$, and $D_{\mathrm{eff}}$, all computed from a combination of data from LYMID and ETS. Note that given the non-negligible contribution to ionization from $\Gamma_{D,\parallel}$ in the SOL, we only calculate $\Gamma_{D}$ and $D_{\mathrm{eff}}$ for $\psi_{n}<$ 1. Table 1 summarizes the range in values for these profiles both at the pedestal top, defined as the point $\Delta$/2 to the left of the center of the mtanh profile, where $\Delta$ is the width of the mtanh (as well as the pedestal). Tabulated as well are their values at the separatrix.
Uncertainties in the profiles in Figure 4 have been computed using standard error propagation, assuming uncorrelated uncertainties from $n_{e}$, $T_{e}$, and $\epsilon_{\mathrm{Ly_{\alpha}}}$. To that error, we add, in quadrature, an error associated with uncertainty in the poloidal alignment of the two diagnostics. For $\epsilon_{\mathrm{Ly_{\alpha}}}$ from LYMID, we assume an error in the magnetic reconstruction, $\Delta R_{\mathrm{EFIT}}$ = 5 mm, and multiply this by the radial derivative of the emissivity profile, $\frac{\partial\epsilon_{\mathrm{Ly_{\alpha}}}}{\partial r}$, computed numerically, i.e. $\Delta\epsilon_{\mathrm{Ly_{\alpha}}}=\Delta R_{\mathrm{EFIT}}\frac{\partial\epsilon_{\mathrm{Ly_{\alpha}}}}{\partial r}$.
For ETS, we use the two-point model [41] to align $n_{e}$ and $T_{e}$ relative to the separatrix, $\psi_{n}=1$. This model states that the power carried by the electrons in the SOL leaves the core entirely at the outer midplane and is transported to the divertor target via conduction only. It uses the Spitzer-Härm conductivity, which is proportional to the parallel temperature gradient in the SOL. Hence, the equality,
can be integrated to find $T_{e}^{\mathrm{sep}}$ at the outer midplane (OMP) [41]. In this equation, $\kappa_{0,e}$ is the Spitzer-Härm conductivity, $R$ is the plasma major radius, $\lambda_{q}$ is the width of the heat flux channel, $B_{\theta}$ is the poloidal magnetic field. In the absence of measurements of the parallel heat flux width, we use the Brunner scaling for $\lambda_{q}$ [42]. We further assume negligible temperature at the divertor (i.e. $T_{e,\mathrm{target}}^{7/2}\ll T_{e,\mathrm{OMP}}^{7/2}$) and estimate the connection length required for the integral as $L_{\parallel}=q_{95}\pi R$, where $q_{95}$ is the safety factor at the 95% flux surface. With $T_{e,\mathrm{OMP}}^{\mathrm{sep}}$ in hand, we find the radial shift required to align the fit to the $T_{e}$ profile at $\psi_{n}=1$ and apply the same shift to the $n_{e}$ before using these shifted kinetic profiles to calculate the rate coefficients used for the neutral inferences.
@@ 3 Density pedestal stiffness
3.1 The role of ionization in setting the pedestal height
We use this methodology to interrogate the role of pedestal ionization in influencing the electron density pedestal height. Figure 5 shows $n_{e}$ plotted against $S_{\mathrm{ion}}$ at the separatrix, $S_{\mathrm{ion}}^{\mathrm{sep}}$. At low values of separatrix ionization, $S_{\mathrm{ion}}^{\mathrm{sep}}\lesssim 15\times 10^{22}$ m-3s-1, $n_{e}$ varies weakly with ionization at the pedestal top and perhaps more strongly at the separatrix. As $S_{\mathrm{ion}}$ continues to grow, however, the trends diverge significantly. At the separatrix, $n_{e}$ increases strongly with $S_{\mathrm{ion}}$, while at the pedestal top there is clear insensitivity of the $n_{e}$ pedestal height to the ionization source. These EDA H-Modes settle at a natural density of $n_{e}^{\mathrm{ped}}\approx 2.2\times 10^{20}$ m-3. Pedestal stiffness on Alcator C-Mod has been observed previously and may be a feature of high density EDAs [18]. In that work, it was shown that varying the neutral source through external gas puffing had little effect on gradient scale lengths in the pedestal as well as the pedestal height. Though similar phenomenology is observed here, a key difference is that the H-modes currently analyzed received no external gas puffing after the L-H transition. Once the H-mode was achieved, pedestal fueling occurred solely through recycled neutrals. Pedestal stiffness in opaque plasmas may thus occur naturally, even in the absence of active gas fueling.
Figure 5: Electron density plotted against volumetric neutral ionization rate at the outer midplane at the separatrix (left) and at the top of the $n_{e}$ pedestal (right). The inset shows an enlarged version of the figure, demonstrating the weak dependence at high $S_{\mathrm{ion}}^{\mathrm{ped}}$ and suggesting a maximum achievable of $n_{e}^{\mathrm{ped}}\approx 2.35\times 10^{20}$ m-3.
@@ 3.2 Flux-gradient relationship in stiff profiles
It has often been assumed that $S_{\mathrm{ion}}$ determines the $n_{e}$ [17, 16]. With this framework, we might expect $\Gamma_{D}$, the integral of $S_{\mathrm{ion}}$, to be proportional to the gradient in the density. In other words, $\Gamma_{D}$ and $\nabla n_{e}$ would be linearly dependent, just as in a classical Fickian diffusive system. Figure 6, however, shows that the experimentally inferred $\Gamma_{D}$ from the Lyα camera, calculated at the top of the pedestal does not grow linearly with $\nabla n_{e}$ at the same location. For $\nabla n_{e}\approx-150\times 10^{20}$ m-4, $\Gamma_{D}$ can vary significantly, anywhere from $6\times 10^{20}$ m-2s-1 to $16\times 10^{20}$ m-2s-1. This is not surprising, given the saturation of $n_{e}^{\mathrm{ped}}$ for vastly different $S_{\mathrm{ion}}$ shown in Figure 5.
@@
Figure 6 indicates that there may be a particular gradient beyond which further ionization becomes ineffective in modifying the density profile, yielding a stiff pedestal. The peak $\nabla n_{e}$ occurs in the mid-range of $\Gamma_{D}$ values, at $\sim$10 ×$\times 10^{20}$ m-2s-1. This is consistent with the rapid growth of $n_{e}^{\mathrm{sep}}$ at fixed $n_{e}^{\mathrm{ped}}$ shown in Figure 5. Discharges at low $\Gamma_{D}^{\mathrm{ped}}$ correspond to those at low $S_{\mathrm{ion}}^{\mathrm{sep}}$ and low $n_{e}^{\mathrm{sep}}$. Those at mid $\Gamma_{D}^{\mathrm{ped}}$ and largest $\nabla n_{e}^{\mathrm{ped}}$ correspond to discharges at $S_{\mathrm{ion}}^{\mathrm{sep}}\approx 15\times$ 1022 m-3s-1 and intermediate $n_{e}^{\mathrm{sep}}$. Finally, the discharges at high $\Gamma_{D}^{\mathrm{ped}}$ but reduced $\nabla n_{e}^{\mathrm{ped}}$ (relative to those at mid $\Gamma_{D}^{\mathrm{ped}}$) correspond to those at highest $S_{\mathrm{ion}}^{\mathrm{sep}}$. These are the discharges with increasing $n_{e}^{\mathrm{sep}}$ but nearly fixed $n_{e}^{\mathrm{ped}}$.
@@
It is evident that these pedestals cannot be fully described with a simple linear diffusive transport model. Since $S_{\mathrm{ion}}$, and thus $\Gamma_{D}$, varies significantly at relatively constant $n_{e}^{\mathrm{ped}}$, it is unlikely that just adding a convective term, proportional to $n_{e}^{\mathrm{ped}}$, to the transport model might explain the large changes to $\Gamma_{D}$ in these pedestals. Instead, it could be that a threshold based model, like that describing heat transport near critical gradients in the core [43], may be necessary for understanding pedestal particle transport. Such a phenomenon in the particle channel has been previously suggested on Alcator C-Mod in the near-SOL for ohmic L-modes [19] and in the pedestal of EDA H-modes undergoing strong puffing [18].
@@
Figure 6: Particle flux at the top of the pedestal plotted against density gradient at the same location. Even at large $\Gamma_{D}$, $-\nabla n_{e}$ does not increase much past $300\times 10^{20}$ m-4. The dashed lines show different slopes, corresponding to different values of $D_{\mathrm{eff}}$. The points lie at a large range of $D_{\mathrm{eff}}$, indicating inconsistency with a purely diffusive transport model having constant $D$.
@@
In the absence of such a model, we opt for using the quotient of the inferred particle flux and the density gradient to obtain an effective particle diffusivity, $D_{\mathrm{eff}}$, calculated using Equation 5. While not fully descriptive of the physics of the edge, $D_{\mathrm{eff}}$ provides a point of comparison for analyzing differences in transport between discharges and as a function of different edge conditions. Experimental inferences of $D_{\mathrm{eff}}$ have been previously made on C-Mod using $\Gamma_{D}$ inferred from the workflow in Section 2.4 [14, 23], although never with in the pedestal with both LYMID and ETS simultaneously. On DIII-D, recent experiments have sought to calculate a diffusive coefficient, $D$ and a convective velocity coefficient, $v$ simultaneously using LLAMA [37]. To do this requires perturbing the edge slightly so as to recover $\frac{\partial n}{\partial t}$ without substantially changing the intrinsic transport properties.
@@
For these stationary discharges, we opt for $D_{\mathrm{eff}}$, and add lines of constant $D_{\mathrm{eff}}$ to Figure 6 to demonstrate the large variation in this proxy for transport in these plasmas. As shown in the rightmost panel of Figure 2 and now apparent from Figures 5 and 6, the rapid growth in $S_{ion}$ and $\Gamma_{D}$ at low $P_{\mathrm{net}}$, as well as the saturation of $n_{e}^{\mathrm{ped}}$, is concurrent with a rapid growth of $D_{\mathrm{eff}}^{\mathrm{ped}}$ below this critical value of $P_{\mathrm{net}}\approx 2.3$ MW.

### v2

@@ 2.4 Inference of neutral density, ionization source, and particle flux
With $n_{e}$, $T_{e}$, and $\epsilon_{\mathrm{Ly}_{\alpha}}$ all mapped to the midplane, we make use of a collisional-radiative model to infer $n_{0}$, $S_{\mathrm{ion}}$, $\Gamma_{D}$, and an effective transport coefficient, $D_{\mathrm{eff}}$. A similar procedure has been employed previously, on C-Mod with Dα measurements [14], and on both C-Mod and DIII-D with Lyα [19, 23, 43, 44, 45]. Equations 3 - 6 summarize the procedure for calculating these quantities.
@@
$\mathcal{P}$ is the photon emissivity coefficient, $S_{H}=\frac{\mathcal{S}}{\mathcal{P}}$ is the ratio between $\mathcal{S}$, the effective ionization coefficient, and $\mathcal{P}$, the photon emissivity coefficient. Neutral and ionized molecular contributions are neglibible for Lyα emission. These expressions highlight the implicit dependence of $S_{\mathrm{ion}}$ (and hence $\Gamma_{D,\perp}$) on $n_{e}$ and $T_{e}$.
Figure 3: Typical kinetic profiles as measured by ETS and a combination of LYMID and ETS measurements. The left panel shows profiles from ETS, with $n_{e}$ (top) and $T_{e}$ (bottom). Shown as open circles are the raw ETS points, including measurement uncertainties. Thick solid lines represent profile fits using the modified hyperbolic tangent fit as described in Section 2.3. Thinner solid lines and the shaded region they envelop represent uncertainties in the fit procedure, as outlined in the same section. The right panel shows those from combined ETS and LYMID analysis. From top to bottom are $n_{0}$, $S_{\mathrm{ion}}$, $\Gamma_{D}$, and $D_{\mathrm{eff}}$. The profiles in purple are at higher collisionality (higher $n_{e}$ and lower $T_{e}$), the relevance of which is explained in Section 4. The ones in green are at lower collisionality (lower $n_{e}$ and higher $T_{e}$). The dotted vertical line in gray denotes the separatrix, located at $\psi_{n}=1.0$. As mentioned in the text, $\Gamma_{D}$ and $D_{\mathrm{eff}}$ are only valid up to the separatrix. The profiles shown here correspond to cases simulated using SOLPS-ITER, described in Section 5.
Table 1: Ranges of values at the location of the top of the $n_{e}$ pedestal and at the separatrix for the discharges analyzed. $n_{e}$ and $T_{e}$ vary substantially at the separatrix, but $T_{e}$ primarily varies at the pedestal top. In contrast, all neutral quantities vary substantially across discharges and at both the pedestal and the separatrix.
@@
$n_{0}$ (1016 m-3) 1.0 – 2.1 1.9 – 4.3
$S_{\mathrm{ion}}$ (1022 m-3s-1) 7.8 – 21 9.3 – 23
$\Gamma_{D}$ (1020 m-2s-1) 4.7 – 18 8.2 – 22
$D_{\mathrm{eff}}$ (m2s-1) 0.025 – 0.13 0.016 – 0.11
@@
Equations 3 and 4 use a number of simplifying assumptions about the relevant atomic processes, which are outlined in [44]. For Equation 5, we ensure a stationary plasma density time window and assume no poloidal asymmetries in $S_{\mathrm{ion}}$, allowing computation of the cross-field particle flux, $\Gamma_{D,\perp}$ by integrating Equation 4 in a slab geometry. For the remainder of the paper, we drop the $\perp$ subscript and express the cross-field particle flux as $\Gamma_{D}$. The assumption that there is negligible poloidal asymmetry in $S_{\mathrm{ion}}$ is likely the strongest simplifying assumption. Using two distinct Lyα views, an in-out poloidal asymmetry has been observed in the ionization source on DIII-D [44, 46]. LYMID does not distinctly view the inboard and outboard side, and an in-out source asymmetry may well be present on C-Mod as well. Regardless, past work has shown that C-Mod is dominated by main-chamber fueling [47, 23]. This lends confidence to using neutral inferences at the midplane to make conclusions about fueling and transport in the C-Mod pedestal. We test this assertion with EIRENE, a 3D neutral simulation code, described in Section 5, and observe good agreement with previous conclusions about main chamber fueling.
@@
Figure 3 shows typical profiles produced with this workflow. The panel on the left shows $n_{e}$ and $T_{e}$ measured by ETS as well as profile fits using Equation 1. The right panel of the figure shows $n_{0}$, $S_{\mathrm{ion}}$, $\Gamma_{D}$, and $D_{\mathrm{eff}}$, all computed from a combination of data from LYMID and ETS. Note that given the non-negligible contribution to ionization from parallel particle flux, $\Gamma_{D,\parallel}$, in the SOL, we only calculate $\Gamma_{D}$ and $D_{\mathrm{eff}}$ for $\psi_{n}\leq$ 1. Table 1 summarizes the range in values for these profiles both at the pedestal top, defined as the point $\Delta$/2 radially inward of the center of the mtanh profile, where $\Delta$ is the width of the mtanh (as well as the pedestal). Only the radial location at the top of the $n_{e}$ pedestal is shown. Tabulated as well are values at the separatrix.
Uncertainties in the profiles in Figure 3 have been computed by perturbing the experimental data points, refitting and recomputing associated quantities multiple times, and then computing statistics on the fits. The perturbations to the experimental points are sampled from a Gaussian distribution with a half-width equal to the experimental uncertainty of the ETS data point. This process is repeated $N$ times and the uncertainty is set to the standard deviation of the ensemble of $N$ perturbed fits. Through sensitivity analysis, it was determined that $N=100$ represented a good compromise between computational time and representation of the uncertainty. This technique enables uncertainty estimation not only in the absolute $n_{e}$ and $T_{e}$ profiles, but also in their gradients, as well as the neutral quantities computed from the plasma profiles. Furthermore, like proper weighted least squares including off-diagonal terms, this uncertainty estimation technique accounts for covariance between fit parameters. The technique is described in greater detail in [45].
As indicated earlier, we do not apply a shift to $\epsilon_{\mathrm{Ly_{\alpha}}}$, as the measurement is made at the midplane and there is no model to constrain the measurement relative to the separatrix. To account for this uncertainty, we compute an error associated with uncertainty in the position of the separatrix as indicated by the magnetic reconstruction computed with EFIT, $\Delta R_{\mathrm{EFIT}}$ = 2 mm. We thus estimate the error in $\epsilon_{\mathrm{Ly_{\alpha}}}$ from LYMID, $\Delta\epsilon_{\mathrm{Ly_{\alpha}}}$, by multiplying $\Delta R_{\mathrm{EFIT}}$ by the radial derivative of the emissivity profile, $\frac{\partial\epsilon_{\mathrm{Ly_{\alpha}}}}{\partial r}$, computed numerically, i.e. $\Delta\epsilon_{\mathrm{Ly_{\alpha}}}=\Delta R_{\mathrm{EFIT}}\frac{\partial\epsilon_{\mathrm{Ly_{\alpha}}}}{\partial r}$. This error is then added in quadrature to the quantities computed from Equations 36.
@@ 3 Density pedestal stiffness
3.1 The role of ionization in the pedestal
We use this methodology to interrogate the role of pedestal ionization in influencing the electron density pedestal height. Figure 4 shows $n_{e}$ plotted against $S_{\mathrm{ion}}$ at the separatrix, $S_{\mathrm{ion}}^{\mathrm{sep}}$. At low values of separatrix ionization, $S_{\mathrm{ion}}^{\mathrm{sep}}\lesssim 15\times 10^{22}$ m-3s-1, $n_{e}$ varies weakly with ionization at the pedestal top and perhaps more strongly at the separatrix. As $S_{\mathrm{ion}}$ continues to grow past $S_{\mathrm{ion}}^{\mathrm{sep}}=15\times 10^{22}$ m-3s-1, however, the trends diverge significantly. At the separatrix, $n_{e}$ continues to increase with $S_{\mathrm{ion}}$, while at the pedestal top there is clear insensitivity of the $n_{e}$ pedestal height to the ionization source. These EDA H-modes settle at a natural density of $n_{e}^{\mathrm{ped}}\approx 2.2\times 10^{20}$ m-3. Pedestal stiffness on Alcator C-Mod has been observed previously and may be a feature of high density EDAs [18]. In that work, it was shown that varying the neutral source through external gas puffing had little effect on gradient scale lengths in the pedestal as well as the pedestal height. Though similar phenomenology is observed here, a key difference is that the H-modes currently analyzed received no external gas puffing after the L-H transition. Once the H-mode was achieved, pedestal fueling occurred solely through recycled neutrals. Pedestal stiffness in opaque plasmas may thus occur naturally, even in the absence of active gas fueling.
Figure 4: Electron density plotted against volumetric neutral ionization rate at the outer midplane at the separatrix (light red squares) and at the top of the $n_{e}$ pedestal (light blue diamonds). While $n_{e}^{\mathrm{sep}}$ continues to increase as $S_{\mathrm{ion}}^{\mathrm{sep}}$ does, $n_{e}^{\mathrm{ped}}$ does not beyond a value of $n_{e}^{\mathrm{ped}}\approx 2.35\times 10^{20}$ m-3.
@@ 3.2 Flux-gradient relationship in stiff profiles
It has often been assumed that $S_{\mathrm{ion}}$ determines $n_{e}$ [17, 16]. With this framework, we might expect $\Gamma_{D}$, the integral of $S_{\mathrm{ion}}$, to be proportional to the gradient in the density. In other words, $\Gamma_{D}$ and $\nabla n_{e}$ would be linearly dependent, just as in a classical Fickian diffusive system. Figure 5, however, shows that the experimentally inferred $\Gamma_{D}$ from the Lyα camera, calculated at the top of the pedestal does not grow linearly with $\nabla n_{e}$ at the same location. For $\nabla n_{e}\approx-150\times 10^{20}$ m-4, $\Gamma_{D}$ can vary significantly, anywhere from $6\times 10^{20}$ m-2s-1 to $16\times 10^{20}$ m-2s-1. This is not surprising, given the saturation of $n_{e}^{\mathrm{ped}}$ for vastly different $S_{\mathrm{ion}}$ shown in Figure 4.
@@
Figure 5 indicates that there may be a particular gradient beyond which further ionization becomes ineffective in modifying the density profile, yielding a stiff pedestal. The peak $\nabla n_{e}$ occurs in the mid-range of $\Gamma_{D}$ values, at $\sim$10 ×$\times 10^{20}$ m-2s-1. This is consistent with the rapid growth of $n_{e}^{\mathrm{sep}}$ at fixed $n_{e}^{\mathrm{ped}}$ shown in Figure 4. Discharges at low $\Gamma_{D}^{\mathrm{ped}}$ correspond to those at low $S_{\mathrm{ion}}^{\mathrm{sep}}$ and low $n_{e}^{\mathrm{sep}}$. Those at mid $\Gamma_{D}^{\mathrm{ped}}$ and largest $\nabla n_{e}^{\mathrm{ped}}$ correspond to discharges at $S_{\mathrm{ion}}^{\mathrm{sep}}\approx 15\times$ 1022 m-3s-1 and intermediate $n_{e}^{\mathrm{sep}}$. Finally, the discharges at high $\Gamma_{D}^{\mathrm{ped}}$ but reduced $\nabla n_{e}^{\mathrm{ped}}$ (relative to those at mid $\Gamma_{D}^{\mathrm{ped}}$) correspond to those at highest $S_{\mathrm{ion}}^{\mathrm{sep}}$. These are the discharges with increasing $n_{e}^{\mathrm{sep}}$ but nearly fixed $n_{e}^{\mathrm{ped}}$. Generally also, discharges at low $S_{\mathrm{ion}}$ and low $\Gamma_{D}$ correspond to those at high $P_{\mathrm{net}}$ and vice versa, as suggested by Figure 2.
@@
It is evident that these pedestals cannot be fully described with a purely diffusive transport model of constant $D$. Since $S_{\mathrm{ion}}$, and thus $\Gamma_{D}$, varies significantly at relatively constant $n_{e}^{\mathrm{ped}}$, it is unlikely that just adding a convective term, proportional to $n_{e}^{\mathrm{ped}}$, to the transport model might explain the large changes to $\Gamma_{D}$ in these pedestals. Instead, it could be that a threshold based model, like that describing heat transport near critical gradients in the core [48], may be necessary for understanding pedestal particle transport. Such a phenomenon in the particle channel has been previously suggested on Alcator C-Mod in the near-SOL for ohmic L-modes [19] and in the pedestal of EDA H-modes undergoing strong puffing [18].
@@
Figure 5: Particle flux at the top of the pedestal plotted against density gradient at the same location. Even at large $\Gamma_{D}$, $-\nabla n_{e}$ does not increase much past $300\times 10^{20}$ m-4. The dashed lines show different slopes, corresponding to different values of $D_{\mathrm{eff}}$. The points lie at a large range of $D_{\mathrm{eff}}$, indicating inconsistency with a diffusive transport model using a constant $D$.
@@
In the absence of such a model, we opt for using the quotient of the inferred particle flux and the density gradient to obtain an effective particle diffusivity, $D_{\mathrm{eff}}$, calculated using Equation 6. While not fully descriptive of the physics of the edge, $D_{\mathrm{eff}}$ provides a point of comparison for analyzing differences in transport between discharges and as a function of different edge conditions. Experimental inferences of $D_{\mathrm{eff}}$ have been previously made on C-Mod using $\Gamma_{D}$ inferred from the workflow in Section 2.4 [14, 23], although never within the pedestal with both LYMID and ETS simultaneously. An important caveat of this steady-state analysis is that in using a single coefficient, one cannot distinguish different types of particle transport, as becomes possible using time-dependent analysis. On DIII-D, for example, recent experiments have sought to calculate both a diffusive coefficient, $D$, and a convective velocity coefficient, $v$, simultaneously using LLAMA [44]. To do this requires perturbing the edge slightly so as to recover $\frac{\partial n}{\partial t}$ without substantially changing the intrinsic transport properties.
@@
For these stationary discharges, we opt for $D_{\mathrm{eff}}$, and add lines of constant $D_{\mathrm{eff}}$ to Figure 5 to demonstrate the large variation in this proxy for transport in these plasmas. As shown in the rightmost panel of Figure 2 and now apparent from Figures 4 and 5, the rapid growth in $S_{\mathrm{ion}}$ and $\Gamma_{D}$ at low $P_{\mathrm{net}}$, as well as the saturation of $n_{e}^{\mathrm{ped}}$, occurs along with a rapid growth of $D_{\mathrm{eff}}^{\mathrm{ped}}$ below the critical value of $P_{\mathrm{net}}\approx 2.3$ MW.